OpenStudy (anonymous):

How many atoms are in 85.1 grams of neon? help ppl I need steps I am so confused

OpenStudy (anonymous):

Given weight of Neon \((\mathsf{Ne})\) = 85.1 grams. Molar mass of Ne = 20 grams. Since no. of moles (n) = \(\cfrac{\textbf{Given weight}}{\textbf{Molar Mass}} \) So, no. of moles (n) of Neon = \(\cfrac{85.1}{20} \) = \(4.255\) Now, use the following formula : \(\textbf{No. of particles in the entity } =\bf{ n \times N_A }\) Where : n = no. of moles \(N_A\) = Avogadro's number = \(\large 6.022 \times 10^{23} \) So, you get : \(\textbf{No. of atoms in Neon} = 4.255 \times 6.022 \times 10^{23}\) Solve it further.

OpenStudy (abb0t):

If you need further clarification, I just explained a step by step process to someone else: http://openstudy.com/study#/updates/51ad25d8e4b06ee3ee25b0c9

OpenStudy (abb0t):

Although I think @Thoughts did a great job explaining.

OpenStudy (anonymous):

Thanks @abb0t . And really, good work there (in the link)